: I drop a tennis ball from a height of h = 5.52 m above an unrealistically huge pile of thick shaving gel. The ball plunges to a depth of d = 18.0 m in the gel before stopping in the gel, at some unknown distance above the floor.

a) What is the magnitude of the ball's acceleration while it's falling through the air? b) Calculate the ball's impact speed v (its speed just before it hits the gel). c) Explain why we should expect the ball's acceleration to be smaller while it's moving through the gel instead of the air; that is, why it should speed up (while falling through the air) more rapidly than it slows down (while stopping in the gel). We're comparing magnitudes only, not taking signs into account. d) Calculate the magnitude a of tennis ball’s stopping acceleration (how rapidly it slowed down in the gel, assuming this occurred at a constant rate). e) How much time t did it take the ball to stop, once it hit the gel? This isn't the time from release to stopping, or from release to impact, but from impact to stopping only.
Answer #1

(A) Assuming the ball is in free fall during its motion through air,

a = - 9.8 m/s2

magnitude , |a| = 9.8 m/s2

--------------------------------

(B) The ball starts from rest. We can use kinematics

v2 - u2 = 2ah

where u is initial velocity

v = sqrt (2ah)

v = sqrt (2 * 9.8 * 5.52)

v = 10.4 m/s

-----------------------------------------------

(C) There is very less air resistance as compared to resistance by the shaving gel because of high viscosity ( this tells us how thick or dense the object is) . Even if we assume that the setup was in vacuum with no air resistance, acceleration will be lot less in gel as compared to in air.

-------------------------------------------------

(D) We know the initial velocity for gel, u = 10.4 m/s

final velocity, v = 0 as ball comes to stop

h = 18 m

so,

v2 - u2 = 2ah

0 - 10.42 = 2 * a * 18

a = - 3 m/s2

question is asking for magnitude

a = 3 m/s2

----------------------------------------------------------------

(E) v = u + at

0 = 10.4 + (-3) t

t = 3.46 sec